OpenStudy (anonymous):

A nurse recorded the heights of the four Aduke children on the chart below. Which child is the tallest? Asha=62 inches Deandre=4 feet 5 inches Juji= 4 1/2 feet malcolm=50 inches

hartnn (hartnn):

do you know how many inches are there in 1 feet ?

OpenStudy (anonymous):

i dont

hartnn (hartnn):

1 feet = 12 inch so how about 4 feet ? how many inches ?

OpenStudy (anonymous):

48

hartnn (hartnn):

yes! so, Deandre=4 feet 5 inches = 48 inches + 5 inches = 53 inches

hartnn (hartnn):

similarly can you find juji's height in inches ?

OpenStudy (anonymous):

approximatly 50 inches

hartnn (hartnn):

4 feet = 48 inches 1/2 feet = 6 inches Juji= 4 1/2 feet = 48+6 = 54 inches got this ?

hartnn (hartnn):

Asha=62 inches Deandre=4 feet 5 inches = 53 inches Juji= 4 1/2 feet = 54 inches malcolm=50 inches now can you tell who is tallest ?

hartnn (hartnn):

which number is greater among 62,53,54,50 ?

OpenStudy (anonymous):

62

hartnn (hartnn):

yes, so Asha is tallest :)

OpenStudy (anonymous):

how

hartnn (hartnn):

because her height is greatest among them

OpenStudy (anonymous):

nvm i got it

hartnn (hartnn):

ok :)